# DK-SLAM: Monocular Visual SLAM with Deep Keypoints Adaptive Learning, Tracking and Loop-Closing

### v1

@@ IV-B Pose Evaluation on the KITTI Dataset
Figure 5: The samples of keypoints detection and matching. Up: matching without two-stage strategy (Ours1). Bottom: matching with two-stage strategy (Ours3).
TABLE III: The ablation study into our DK-SLAM modules.
Method MAML Two-Stage Metric 00 02 05 07 09 10 Avg
Ours1 $\checkmark$ $t_{rel}$ 2.88 - 2.80 1.76 2.70 - -
$r_{rel}$ 0.41 - 0.35 0.64 0.24 - -
Ours2 $\checkmark$ $t_{rel}$ 2.92 4.42 2.48 2.11 4.03 8.48 4.07
$r_{rel}$ 0.38 0.34 0.34 0.46 0.20 0.23 0.325
Ours3 $\checkmark$ $\checkmark$ $t_{rel}$ 2.57 4.38 1.98 1.11 2.97 7.03 3.34
$r_{rel}$ 0.31 0.27 0.26 0.28 0.29 0.23 0.27
IV-C Pose Evaluation on the EuRoC Dataset
We validate our proposed DK-SLAM on the EuRoC dataset, addressing scale ambiguity with umeyama alignment [37] for monocular vision methods. Comparisons include ORB-SLAM3, DSO, DSM [38], SVO, and LIFE-SLAM [8]. Results, derived from [6] for ORB-SLAM3, DSO, DSM, and SVO, and [8] for LIFE-SLAM, utilize absolute translation error as the experiment metric . The proposed SLAM exhibits exceptional positioning performance, leveraging a two-stage tracking strategy for precise feature point location and reduced incorrect matching. Learnt features demonstrate robust detection in low-light environments, fostering stable feature tracking. The online learning-based deep BoW excels in loop detection, evident in MH02 where the proposed SLAM outperforms ORB-SLAM3. This advantage is attributed to learned features constructing BoW, providing detailed scene distinction, and online training for swift adaptation to varying environments.
IV-D Ablation Study
Table III summarizes the results of our ablation study, assessing the impact of our MAML-based deep feature training and coarse-to-fine feature tracking. In ”Ours1,” lacking a coarse-to-fine tracking strategy, reliance on a constant velocity motion model for keypoint matching leads to errors, resulting in inaccurate correspondence and tracking failure. Both ”Ours1” and ”Ours3” use a keypoint search radius of 7.
Figure 5 shows a failure case for ”Ours1” in accurate feature matching due to uniform motion model challenges. In contrast, ”Ours3” uses a two-stage strategy for stable tracking, combining a semi-direct method for coarse pose estimation and refined feature matching. ”Ours2” with the original Superpoint strategy performs slightly worse than ”Ours3” trained with MAML. In ”Ours3,” MAML is applied to train the local feature extractor by iteratively enhancing generalization on support and query sets, resulting in a robust and adaptable detector. The enhanced SLAM performance attributed to MAML is primarily observed in feature matching, as depicted in Figure 6. Using Superpoints for matching increases the number of matches, addressing ORB-SLAM3’s limitations in texture-less areas. ORB-SLAM3’s descriptors, describing lighting changes around corners, are susceptible to environmental variations. In contrast, learning-based local feature methods like Superpoint capture robust deep features unaffected by lighting changes, ensuring stable tracking across consecutive frames.
@@ IV-D Ablation Study
TABLE IV: The average number of matching points for ORB-SLAM3, and DK-SLAM without (Ours2) or with MAML feature training (Ours3).
@@
Figure 6: The samples of keypoints detection and matching. From top to bottom: ORB-SLAM3 matching (ORB3), DK-SLAM without (Ours2) or with MAML feature training (Ours3).
@@
We perform ablation studies on the proposed loop closure method to assess its performance. Precision-Recall metrics on KITTI 00, 05, and 06 sequences are illustrated in Figure 7, with additional evaluation at 100% precision. The quantitative results, presented in Table V, reveal our deep BoW’s superior recall rate compared to traditional BoWs. Notably, in the KITTI 06 sequence, our BoW achieves a remarkable 100% recall rate. In contrast, the recall rate of the iBoW, belonging to the traditional BoW category, is significantly lower, suggesting that handcrafted descriptors might struggle to accurately identify loop nodes in a sequence. The learned local descriptor captures high-level information, enhancing robustness across diverse environments. This feature stability, unaffected by lighting changes, results in superior loop detection performance.
@@ V Conclusion
This work presents DK-SLAM, a monocular visual SLAM with adaptive learning features, featuring a learned local feature detector, a coarse-to-fine matching strategy, and an online binary learned feature BoW. To enhance generalization, we employ MAML for adaptive training, yielding an extractor with robust capabilities. Coarse relative poses between frames are estimated via a semi-direct method for accurate feature point matching. An online binary learned feature BoW corrects SLAM accumulation errors. Our proposed DK-SLAM outperforms representative baselines, e.g. ORB-SLAM3, on the KITTI and EuRoC datasets. However, the GPU-based online front-end faces efficiency challenges during information transmission to the CPU-based back-end. Future work will explore knowledge distillation for parameter compression and SLAM framework modifications to enhance efficiency.

### v2

@@ IV-C Pose Evaluation in the UAV Scenario
Figure 7: The samples of keypoint detection and matching. Up: matching without two-stage strategy (Ours1). Bottom: matching with two-stage strategy (Ours3).
@@
TABLE IV: The average number of matching points for ORB-SLAM3, and DK-SLAM without (Ours2) or with feature meta learning (Ours3).
@@
Figure 8: The samples of keypoints detection and matching. From top to bottom: ORB-SLAM3 matching (ORB3), DK-SLAM without (Ours2) or with keypoint meta learning (Ours3).
IV-D Ablation Study
Table III summarizes the results of the ablation study into the key modules within DK-SLAM system, assessing the impact of our MAML-based deep keypoint meta learning and coarse-to-fine keypoint tracking. For a fair comparison, "Ours1", "Ours2" and "Ours3" all use a keypoint search radius of 7.
1) The analysis of keypoint meta learning module: the module’s effectiveness is evident when comparing "Ours2" and "Ours3." "Ours2," which employs the original SuperPoint strategy without our keypoint meta-learning approach, performs slightly worse than "Ours3," which integrates MAML-based keypoint meta-learning. In "Ours3," MAML is applied to train the local feature extractor by iteratively enhancing generalization on support and query sets, resulting in a robust and adaptable detector. The batch size of "Ours2" and "Ours3" is set as 8. Both models were trained 150K iterations. In "Ours3", we split every training batch to 4 mini-batch support set and 4 mini-batch query set. As mentioned in algorithm 1, the number of update iterations in inner loop is 4.
2) The analysis of coarse-to-fine tracking module: in "Ours1," lacking a coarse-to-fine tracking strategy, reliance on a constant velocity motion model for keypoint matching leads to errors, resulting in inaccurate correspondence and tracking failure. Figure 7 shows a failure case for "Ours1" in accurate feature matching due to uniform motion model challenges. In contrast, "Ours3" uses a two-stage strategy for stable tracking, combining a semi-direct method for coarse pose estimation and refined feature matching.
The enhanced SLAM performance attributed to proposed keypoint meta learning and two-stage matching strategy can also be observed in feature matching results, as depicted in Figure 8. Using SuperPoint for matching increases the number of matches, addressing ORB-SLAM3’s limitations in texture-less areas. ORB-SLAM3’s descriptors, describing lighting changes around corners, are susceptible to environmental variations. In contrast, the original learning-based local feature methods such as SuperPoint capture robust deep features unaffected by lighting changes, ensuring stable tracking across consecutive frames.
@@ IV-D Ablation Study
3) The analysis of loop-closing module: we further perform ablation studies on the proposed loop closure method to assess its performance. Precision-Recall metrics on KITTI 00, 05, and 06 sequences are illustrated in Figure 9, with additional evaluation at 100% precision. The quantitative results, presented in Table V, reveal our deep BoW’s superior recall rate compared to traditional BoWs. Notably, in the KITTI 06 sequence, our BoW achieves a remarkable 100% recall rate. In contrast, the recall rate of the iBoW, belonging to the traditional BoW category, is significantly lower, suggesting that handcrafted descriptors might struggle to accurately identify loop nodes in a sequence. The learned local descriptor captures high-level information, enhancing robustness across diverse environments. This feature stability, unaffected by lighting changes, results in superior loop detection performance.
@@ V Conclusion
This work presents DK-SLAM, a monocular visual SLAM with deep keypoint meta learning, a coarse-to-fine matching strategy, and an online binary learned feature BoW. To enhance generalization, we employ MAML for local feature training, yielding a feature extractor with robust capabilities in unseen scenarios. Coarse relative poses between frames are estimated via a semi-direct method for accurate feature point matching. An online binary learned feature BoW corrects SLAM accumulation errors. Our proposed DK-SLAM outperforms representative baselines, e.g. ORB-SLAM3, on the KITTI and EuRoC datasets. However, the GPU-based online front-end faces efficiency challenges during information transmission to the CPU-based back-end. Future work will explore knowledge distillation for neural network parameters compression and further SLAM framework modifications to enhance efficiency.